# Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation

### v3

@@ IV-A Case description
The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is kW/kVar. The step sizes are chosen with a trial-and-error strategy [24] as $\alpha^{d}=10^{6}$, $\alpha=5\times 10^{-4}$, and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 0.6%. The primal and dual regularization is used to establish the theoretical convergence result. The impact of some small regularization parameters on numerical results is negligible. In the simulation, we set them as $10^{-4}$. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (17), i.e. we assume that every setpoint in Algorithm 1 is implemented for 1 second. Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient. For a series implementation of a 4-hour simulation, i.e. in total 14400 iterations, it takes 8.1 seconds, averaging 0.56 milliseconds per iteration.
@@
IV-B Static case results
@@
In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach. The simulation uses data at 12:00. Figure 6 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 200 total iterations, i.e. 33 outer OFO iterations. The approach also utilizes the available voltage limit as efficiently as the centralized one. Figure 7 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (6).
IV-C Dynamic case results
@@
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizers of (6) without waiting for them to converge. Figure 8a shows significant voltage limit violations without control. Figures 8b and 8c demonstrate that the centralized and our proposed distributed nested feedback optimization approaches both successfully enforce the voltage limit, although temporary voltage limit violations are inevitable due to the corrective nature of these algorithms. The simple two-metric approach and the heuristic approach in [23] result in prolonged voltage limit violations as seen in Figs. 8d and 8e, respectively. Figure 8f shows that the distributed approach in [24] does not track the time-varying optimizers well and experiences oscillations at the end of the simulation period. The discrepancy between the actual setpoints and algorithm iterates (without projection) potentially renders the algorithm slow and unstable. Figure 8g shows that the local droop control approach does not enforce the voltage limit due to lack of coordination. Finally, for clarity, Fig. 8h presents voltages for the most remote, and therefore, the most sensitive bus (bus 95) under different approaches.
@@
Furthermore, Fig. 9 shows reactive power profiles under different approaches. Shown in Figs. 9a and 9b, our proposed distributed approach achieves similar profiles with the centralized one, with an average setpoint deviation of $-1.3\times 10^{-3}$ kVar. The two-metric approach and the heuristic approach do not work well in this case study. The approach in [24] experiences oscillations. The local droop control approach does not in theory or practice track the optimizers of (6).
@@
TABLE I: Average voltage violations for various approaches.
Approach AVV (pu) w.r.t. centralized
No control $1.5\times 10^{-2}$ 192x
Centralized $8.0\times 10^{-5}$ 1x
Proposed $1.6\times 10^{-4}$ 2x
Two-metric $2.5\times 10^{-3}$ 31x
Tang 2024 [23] $3.3\times 10^{-3}$ 41x
Qu 2020 [24] $8.6\times 10^{-4}$ 11x
Droop control [37, 38] $4.5\times 10^{-4}$ 6x
@@
Table I shows the AVV values for bus 95 (the most sensitive bus) under various approaches. An AVV of $1.6\times 10^{-4}$ pu is achieved for the proposed approach, which is only higher than that of the centralized approach. The latter however requires a centralized communication architecture and is thus susceptible to single-point failures. The proposed nested feedback approach significantly outperforms other existing distributed or local approaches.
@@ V Conclusion
In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach. We theoretically analyzed its convergence. Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed existing distributed and local approaches.
The proposed distributed approach does not extend to congestion management while the centralized does so. The current presentation is based on radial systems but the approach can be readily extended to meshed systems using the approach in [3] to construct the sensitivity matrix $\mathbf{X}$. Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix. Finally, the approach works for either active or reactive power control, but not joint control. This is because the observation in Section III-B does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.

### v4

@@ IV-A Case description
(a) Trajectories of the cost function
(b) Distribution of the relative cost function (w.r.t. the optimum)
Figure 5: Results of 1000 random problem instances.
@@
The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is kW/kVar. For the inner loop, $\alpha^{u}$ is chosen as $0.99\cdot 2/\lambda_{max}(\mathbf{X})$. For the outer loop, the step sizes are chosen with a trial-and-error strategy [28] as $\alpha^{d}=10^{6}$ and $\alpha=3\times 10^{-4}$. Following the guideline in Remark 2, the regularization parameters $r_{p}$ and $r_{d}$ are set to $10^{-4}$. The algorithm performs well with such small regularization, likely due to the strongly convex objective function in (6). For each generation and load data point, 3 OFO iterations are run which yields tentative reactive power setpoints, followed by 10 inner iterations, i.e. we assume that each inner iteration takes 200 milliseconds. It is crucial that the system responds faster than the DER setpoint update rate. In our application, setpoints are updated on a (sub-)second timescale, and the system is assumed to react even more quickly. Similarly, [40, 1, 2] consider DER update time of 0.1, 0.33, and 1 second, respectively. Another relevant example is droop control, where voltage measurements and setpoint updates occur in real time, typically within the subsecond range [41]. Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient. For a series implementation of a 4-hour simulation, i.e. in total 7200 iterations, it takes 6.7 seconds, averaging 0.93 milliseconds per iteration.
@@
Figure 8: Loss comparison between the objectives of reactive power minimization and loss minimization.
@@
IV-B Convergence of inner loop
To demonstrate the convergence of the inner loop, 1000 random problem instances for (13) are generated. Figure 5a illustrates that the cost function drops significantly within a few iterations, leading to rapid algorithm convergence. Additionally, Fig. 5b depicts the distribution of the relative cost function with respect to the optimum. After approximately 10 iterations, the cost function stabilizes. Thus, in this study, we set the inner loop to run for 10 iterations, ensuring a balance between solution accuracy of the inner loop and the update frequency of the outer loop.
IV-C Static case results
@@
In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach. The simulation uses data at 12:00. Figure 6 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 60 outer iterations. The approach also utilizes the available voltage limit as efficiently as the centralized one. Figure 7 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (6). Finally, Fig. 8 demonstrates that under overvoltage issues caused by excessive PV generation, the chosen objective function of minimizing reactive power usage aligns with minimizing system losses. However, under undervoltage conditions, the outcome may differ, as increased reactive power injection could potentially help lower system losses. This highlights a tradeoff between the distributed communication architecture and the flexibility in selecting optimization objectives.
IV-D Dynamic case results
@@
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizer of (6) without waiting for them to converge. Figure 9a shows significant voltage limit violations without control. Figures 9b and 9c demonstrate that the centralized and our proposed distributed nested feedback optimization approaches both successfully enforce the voltage limit, although temporary voltage limit violations are inevitable due to the corrective nature of these algorithms. The simple two-metric approach and the heuristic approach in [27] result in prolonged voltage limit violations as seen in Figs. 9d and 9e, respectively. Figure 9f shows that the distributed approach in [28] does not track the time-varying optimizers well and experiences oscillations at the end of the simulation period. The discrepancy between the actual setpoints and the algorithm iterates (without projection) potentially renders the algorithm slow and unstable. Figure 9g shows that the local droop control approach does not enforce the voltage limit due to lack of coordination. Finally, for clarity, Fig. 9h presents voltages for the most remote, and therefore, the most sensitive bus (bus 95) under different approaches.
@@
Furthermore, Fig. 10 shows reactive power profiles under different approaches. Shown in Figs. 10a and 10b, our proposed distributed approach achieves similar profiles with the centralized one, with an average setpoint deviation of 0.01 kVar. The two-metric approach and the heuristic approach do not work well in this case study. The approach in [28] experiences oscillations. The local droop control approach does not in theory or in practice track the optimizer of (6).
@@
TABLE I: Average voltage violations and losses for various approaches.
Approach AVV (pu) AVV w.r.t. central. Loss (kWh)
No control $1.5\times 10^{-2}$ 192x 48.2
Centralized $8.0\times 10^{-5}$ 1x 64.0
Proposed $9.9\times 10^{-5}$ 1.2x 64.0
Two-metric $2.5\times 10^{-3}$ 31x 59.3
Tang 2024 [27] $3.3\times 10^{-3}$ 41x 57.4
Qu 2020 [28] $8.6\times 10^{-4}$ 11x 65.9
Droop control [42] $4.5\times 10^{-4}$ 6x 67.8
@@
Table I shows the AVV values for bus 95 (the most sensitive bus) and system losses under various approaches. An AVV of $9.9\times 10^{-5}$ pu is achieved for the proposed approach, which is slightly higher than that of the centralized approach. The latter however requires a centralized communication architecture and is thus susceptible to failure of the central coordination unit. The proposed nested feedback approach significantly outperforms other existing distributed or local approaches. Because reactive power is absorbed to mitigate voltage rise, system losses increase. The proposed approach results in the same losses as the centralized method.
IV-E Impact of measurement noise
Figure 11: Voltages for bus 95 under measurement noise.
This section examines the impact of measurement noise on the voltage regulation performance. The noise is modeled using a zero-mean Gaussian distribution with a standard deviation of $\sigma$. Figure 11 illustrates its effect. As measurement noise increases, voltage fluctuations become more pronounced. Nevertheless, the controller successfully maintains voltages within safe limits. The temporary voltage limit violations can be mitigated by setting a tighter limit, e.g. 1.048 pu.
IV-F Impact of node failures
Figure 12: Voltages for bus 95 under a failed node.
Figure 13: Impact of node failres on AVV.
In this section, we demonstrate the robustness of the proposed strategy against random node failures. When a node fails, its neighbouring nodes are assumed to bypass it and continue communication. Figure 12 illustrates an example voltage profile with a failed node, showing that the system maintains the voltage limit despite a random node failure. Figure 13 depicts how AVV increases with the number of failed nodes but remains within a reasonable range, indicating the algorithm’s robustness. Unlike centralized approaches, where failure of the central coordination unit leads to a complete loss of voltage regulation, the proposed method continues to function effectively. Enhancing system resilience through the development of failure detection and bypass mechanisms is an important direction for future research.
IV-G Impact of update frequency
Figure 14: AVV under various update frequency.
Figure 14 quantifies the impact of the OFO controller update frequency on its voltage regulation performance. Note that the simulated generation and load data have a resolution of 6 seconds. As expected, more frequent updates improve performance. However, even with an outer iteration length of 24 seconds, the performance deterioration remains limited. This suggests that the method retains effectiveness even under slower update rates. While we assume relatively fast update in the inner loop (e.g. 200 milliseconds per iteration), we acknowledge that this may be optimistic in large or deep networks. Therefore, this section serves to assess the controller’s robustness to slower updates, which relaxes the need for low-latency communication. The optimal update frequency depends on the communication infrastructure and requires a cost-effectiveness analysis. Future work will focus on analyzing the scalability of the controller under more realistic communication constraints, including increased latencies and asynchronous operation.
Figure 15: Voltages for bus 95 under various objective functions.
IV-H Impact of the objective function
In this section, we compare our objective function with the alternative objective of minimizing $\mathbf{q}^{\intercal}\mathbf{X}\mathbf{q}$, as adopted in [3, 12, 20]. Figure 15 shows that this weighted objective function (i.e. weighted by $\mathbf{X}$) also effectively enforces the voltage limit in the dynamic case. The voltage profile obtained using the weighted cost function yields a lower average voltage violation (AVV) of $3.0\times 10^{-5}$ pu but exhibits greater variation. Given that the system faces overvoltage issues, the voltage constraints are expected to be binding at the optimal solution. Accordingly, we use 1.05 pu as the reference voltage for the most remote bus. Under this setting, the voltage profile from the weighted cost function results in an average deviation of $6.3\times 10^{-4}$ pu, which is significantly higher than that achieved with our proposed formulation ($2.3\times 10^{-4}$ pu). Moreover, the weighted formulation assigns higher costs to PV inverters located at the end of distribution feeders, leading to suboptimal and unfair utilization of available reactive power capacity. Consequently, it results in a total reactive power use of 460.4 kVarh, compared to 390.9 kVarh in our formulation. Reducing reactive power absorption can be beneficial for lowering system losses and extending PV inverter lifetime. In this simulation, the weighted cost function yields a loss of 65.4 kWh, which is higher than the 64.0 kWh loss achieved with our formulation. Finally, our formulation is flexible and can accommodate this alternative objective function if considered beneficial.
IV-I Active power control
Figure 16: Voltages for bus 95 under various objective functions with active power control.
To further demonstrate the significance of the proposed formulation, we apply the algorithm for active power control to mitigate voltage limit violations in scenarios where reactive power capacity is insufficient. As shown in Fig. 16, both formulations successfully enforce the voltage limit. However, the weighted formulation leads to 17.5% active power curtailment (1649.7 kWh generated out of a possible 2000.7 kWh), whereas the proposed formulation reduces curtailment to only 7.7% (1847.2 kWh generated). This highlights the practical advantage of the proposed approach.
@@ V Conclusion
In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach. We theoretically analyzed its convergence. Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed existing distributed and local approaches. The proposed distributed approach does not extend to congestion management while the centralized does so. Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix. Finally, the approach can be extended to joint active and reactive power control.